OpenStudy (anonymous):

if z is a complex number, find all z that satisfy z^2 + 2z + 1 =0

OpenStudy (slaaibak):

\[z = -1 + 0i\]

OpenStudy (anonymous):

can you really just factor normally?

OpenStudy (slaaibak):

yes, it's (z+1)^2=0

OpenStudy (anonymous):

that makes sense to me, thanks

OpenStudy (slaaibak):

or use this http://www.purplemath.com/modules/quadform.htm

OpenStudy (anonymous):

how about if the equation is z^2 + 2(conjugate z) + 1 = 0

OpenStudy (anonymous):

zarkon are you there?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

can you help me with this one? if z is a complex number, find all z that satisfy z^2 + 2(conjugate z) + 1 = 0

OpenStudy (zarkon):

let \(z=a+bi\) plug into your equation set the real and imaginary parts equal to zero...solve for a,b

OpenStudy (anonymous):

how can I solve for a and b when I only have 1 equation?

OpenStudy (zarkon):

I get \(z=-1,1\pm2i\)

OpenStudy (zarkon):

\(z^2+2\overline{z}+1=0\) \((a+bi)^2+2(a-bi)+1=0\) expand...you will and imaginary part and a real part (thus you will have two equations) solve for a,b

OpenStudy (zarkon):

\(a^2+2a-b^2+1+(2ab-2b)i=0\) thus \(a^2+2a-b^2+1=0\) and \(2ab-2b=0\)

OpenStudy (anonymous):

so in order to have that complex thing be 0 either the real or the imaginary part has to be 0 which means 2 equations that you can solve. Brilliant!

OpenStudy (zarkon):

then both have to be zero a+bi=0 same as a+bi=0+0i equate real and imaginary parts

OpenStudy (anonymous):

right woops

OpenStudy (anonymous):

that makes a lot of sense though. Thanks for your help!

OpenStudy (zarkon):

no problem